OpenStudy (anonymous):

help please !!! y=x^2+5x+6 I need to know the vertex and ordered pair??? Don't understand the formula. Is the vertex -2.5, 4.75? And what is the ordered pair?

OpenStudy (phi):

the x value of the vertex is -b/2a where a and b are the numbers in a x^2 + bx + c in this case a is 1 (x^2 is the same as 1* x^2 , so a is 1) b is +5 so the x value of the vertex is - 5/2 use that value for x in the equation to find y y=x^2+5x+6 with x= -5/2 (or -2.5) y= (-5/2)^2 + 5* -5/2 + 6 can you do the arithmetic to find the y value of the vertex ?

OpenStudy (anonymous):

i dont know how to

OpenStudy (phi):

is your calculator broken? you can type that into google and it will figure it out